OpenStudy (anonymous):

A wine maker has two vats of wine, one with 5% alcohol content and the other with a 20% alcohol content. He wishes to blend them to produce 1L of wine with a 14% alcohol content. How many milliliters of each are required? Show how to use a system of equations to solve this problem.

OpenStudy (anonymous):

hmmm what if instead of trying so hard to get 14 he and a friend sit there and test it until theyre content with the alcohol level. (sorry i just needed one point to be in 50 i know this doesnt help.)

OpenStudy (anonymous):

wow thanks

OpenStudy (mertsj):

Let x = amount of 5% wine to use. Let y = amount of 20% wine to use x+y=1 .05x+.20y=.14(1)

OpenStudy (anonymous):

then just solve for x and y?

OpenStudy (mertsj):

Yes

OpenStudy (anonymous):

since I am looking for milliliters though would I change the one to 1000?

OpenStudy (mertsj):

It would probably be easiest to multiply the second equation by 100 to get: 5x+20y=14 Then multiply the first equation by -5 and add the two together.

OpenStudy (mertsj):

You could change to ml at the end.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so I have the systems x+y=1 and 5x+20y=14 and now I solve for y??

OpenStudy (mertsj):

Yes. Those are the two equations in your system. Solve for both x and y because the problem asks how many mL of each.

OpenStudy (anonymous):

ok so I get 4x-19y=13 to start?

OpenStudy (mertsj):